OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x4 + 36x2 f(x) = x2(6x + i)2 f(x) = x2(x + 6i)2 f(x) = x2(6x + i)(6x - i) f(x) = x2(x + 6i)(x - 6i)

OpenStudy (cwrw238):

= x^2(x^2 + 36) = x^2(x + 61)(x - 6i)

OpenStudy (cwrw238):

thats x^2(x + 6i)(x - 6i)

OpenStudy (anonymous):

Thank you so much! just needed clarification

OpenStudy (cwrw238):

yw